The length of a rectangle is 5cm longer than the width. The perimiter is 54 cm. Find the demensions of the rectangle?
Answer by oberobic(2304)

L = W+5
2(L+W) =54
substitute
2((W+5) +W) = 54
W+5+W = 27
2W = 22
W = 11
L = W+5
L = 16
.
Check perimeter
.
2(L+W) = 2(27) = 54
Correct.
.
Answer: Length = 16, Width = 11
.
Done.